Circular No. 6908

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 K1 (MUELLER) Jean Mueller reports her discovery of a comet in the course of the

second Palomar Sky Survey. The following positions are available:

1998 UT  α_{2000}  δ_{2000}  m_{1} Observer
May 16.20104 10^{h}38^{m}17.^{s}53 +32^{o}34′25′′.6 17.5 Mueller
17.32014 10 38 40.0 +32 18 36  Gal
17.91916 10 38 50.97 +32 09 44.0 16  Tichá
17.93014 10 38 51.16 +32 09 35.1  _{”}
18.15106 10 38 55.66 +32 06 24.6 16.4 Hergenrother
18.27426 10 38 58.08 +32 04 37.7  _{”}

J. Mueller (Palomar). 1.2-m Oschin Schmidt. Some condensation and a very short tail to the east. Comet near corner of 55-min exposure. Measurer Mueller. Reduction by G. V. Williams.

R. Gal and V. Hradecky (Palomar). 1.5-m reflector + CCD. Object diffuse

with tail to the east. Position deduced as offset from field center. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. Meas-

urer Tichý. C. W. Hergenrother (Mt. Hopkins). 1.2-m reflector + CCD.

The following parabolic orbital elements, by the undersigned, are quite uncertain:

T = 1998 May 23.511 TT  ω = 145.^{o}716
Ω = 14.925 }2000.0
q = 3.73200 AU  i = 37.770
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 17 10^{h}38.^{m}55 +32^{o}23.^{′}0 3.549 3.732 92.^{o}5 15.^{o}7 16.0
22 10 40.35 +31 11.0 3.612 3.732 88.9 15.7 16.0
27 10 42.52 +29 58.8 3.677 3.732 85.3 15.7 16.0
June 1 10 45.02 +28 46.5 3.742 3.733 81.7 15.6 16.1
6 10 47.83 +27 34.3 3.808 3.734 78.2 15.4 16.1
11 10 50.91 +26 22.3 3.874 3.736 74.7 15.2 16.2
16 10 54.23 +25 10.5 3.939 3.738 71.2 14.9 16.2
21 10 57.78 +23 59.1 4.005 3.741 67.8 14.6 16.2
26 11 01.54 +22 48.1 4.070 3.744 64.4 14.2 16.3
1998 May 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden